# Disparate Impact on Group Accuracy of Linearization for Private Inference

### v1

@@ Acknowledgements
This research is partly funded by NSF grants SaTC-2345483, RI-2334936, and CAREER-2401285. F. Fioretto is also supported by a Google Research Scholar Award and an Amazon Research Award. The views and conclusions of this work are those of the authors only.
@@ C.1 Settings
For SNL, we train each base model for 160 epochs, and then perform the SNL finetuning step after ReLU linearization. These values are present in the official implementation of this algorithm provided by the authors of Cho et al. [2022], and we use this implementation off-the-shelf.
For DeepReDuce, we use the official implementation provided by the authors of Jha et al. [2021] off-the-shelf which trains each model for 200 epochs.
C.2 For Accuracy
@@ C.2.1 Mitigation
C.3 Mitigation: Gradient Norms and Distance to the Decision Boundary

### v3

@@ 7 Discussion
Limitations. Our theoretical framework proposes a bound for the residual loss of a DNN model and its linearized counterpart, elucidating their relationship with group norms, group Hessians, and the norm of the weight distance between the two models’ parameters. While our analysis introduces proxies to articulate the connection between group size, linearized ReLU proportion, and accuracy drop, examining the actual relationship between these quantities remains an open question. Addressing this gap presents an avenue for future research, with an opportunity to formalize fundamental limits of disparate impact mitigation strategies through non-trivial lower bounds on loss differences. We note that the proposed mitigation strategy necessitates access to protected group information. Although this information is frequently available, the requirement may restrict the applicability of the proposed mitigation in contexts where such information may be missing. To address this limitation, future research could explore the integration of our mitigation approach with tailored loss functions to promote fairness without relying on demographic information (cf. Lahoti et al. (2020)) or with a context of distributionally robust objectives (cf. Hashimoto et al. (2018)).
@@ Acknowledgements
This research is partially supported by NSF grants 2133169, 2232054 and NSF CAREER Award 2143706. Fioretto is also supported by an Amazon Research Award and a Google Research Scholar Award. Its views and conclusions are those of the authors only.
Impact Statement
Our contributions shed light on the Matthew effect within accuracy reduction caused by ReLU linearization, offering a mathematical intuition of the observed disparities, demonstrating the widespread presence of these effects across algorithms and models, and proposing a viable mitigation strategy. The proposed strategy not only addresses fairness concerns but also manages to preserve overall accuracy to a commendable extent.
The evidence presented, particularly the stark accuracy disparities illustrated in Figure 1, calls for a reevaluation of model optimization strategies, especially in the use of datasets reflecting real-world diversity. The findings emphasize the necessity for approaches that consider the differential impact of linearization on various subgroups, advocating for a more inclusive and fair machine learning ecosystem.
@@ C.1 Settings
For SNL, we train each base model for 160 epochs, and then perform the SNL finetuning step after ReLU linearization. These values are present in the official implementation of this algorithm provided by the authors of Cho et al. (2022), and we use this implementation off-the-shelf.
For DeepReDuce, we use the official implementation provided by the authors of Jha et al. (2021) off-the-shelf which trains each model for 200 epochs.
C.2 Accuracy drop analysis
@@ Appendix D Mitigation
D.1 Mitigation: Gradient Norms and Distance to the Decision Boundary
@@
D.2 Mitigation: an ablation study of the parameter $\mu$
(a) SNL on UTKFace race ResNet18 mitigation with $\mu=0.0005$
(b) SNL on UTKFace race ResNet18 mitigation with $\mu=0.005$
(c) SNL on UTKFace race ResNet18 mitigation with $\mu=0.01$
Figure 15: We report the plots for the ablation study of the parameter $\mu$ for the SNL on UTKFace (race). We consider $\mu\in\{0.0005,0.005,0.01\}$. In general we observe that these values of $\mu$ do not result in significant changes for the performance of the mitigation strategy. For the task at end, we could then say that there exists a neighborhood of values for $\mu$ where our mitigation consistently benefits the considered linearization techniques.